# Adversarial Federated Consensus Learning for Surface Defect Classification Under Data Heterogeneity in IIoT

### v1

@@ Abstract
The challenge of data scarcity hinders the application of deep learning in industrial surface defect classification (SDC), as it’s difficult to collect and centralize sufficient training data from various entities in Industrial Internet of Things (IIoT) due to privacy concerns. Federated learning (FL) provides a solution by enabling collaborative global model training across clients while maintaining privacy. However, performance may suffer due to data heterogeneity—discrepancies in data distributions among clients. In this paper, we propose a novel personalized FL (PFL) approach, named Adversarial Federated Consensus Learning (AFedCL), for the challenge of data heterogeneity across different clients in SDC. First, we develop a dynamic consensus construction strategy to mitigate the performance degradation caused by data heterogeneity. Through adversarial training, local models from different clients utilize the global model as a bridge to achieve distribution alignment, alleviating the problem of global knowledge forgetting. Complementing this strategy, we propose a consensus-aware aggregation mechanism. It assigns aggregation weights to different clients based on their efficacy in global knowledge learning, thereby enhancing the global model’s generalization capabilities. Finally, we design an adaptive feature fusion module to further enhance global knowledge utilization efficiency. Personalized fusion weights are gradually adjusted for each client to optimally balance global and local features, tailored to their individual global knowledge learning efficacy. Compared with state-of-the-art FL methods like FedALA, the proposed AFedCL method achieves an accuracy increase of up to 5.67% on three SDC datasets.
@@ I Introduction
In recent years, the integration of artificial intelligence with IIoT has provided new automation solutions across various industries, significantly enhancing efficiency and effectiveness[7, 8, 9, 10, 11, 12, 13, 14]. Traditional machine learning classifiers, such as support vector machine (SVM) and random forest [7, 8, 9], have been utilized for automated SDC. These classifiers are typically coupled with some shape-based and texture-based feature extraction methods [10], such as Fourier descriptor and local binary pattern. However, they suffer from limited feature extraction capabilities due to reliance on manual features, inadequately capturing the diversity of surface defects, thus limiting generalization and adaptability in SDC tasks. On the contrary, deep learning (DL)-based methods are capable of learning abstract representations autonomously, without the need for manual feature engineering. In particular, DL-based classifiers built on convolutional neural network (CNN) architectures, such as ResNet and MobileNet, consistently demonstrate superior performance in SDC tasks [11, 12, 13].
Nonetheless, DL-based SDC methods require sufficient data to develop robust models, which a single entity often cannot provide [15]. To overcome this limitation, data must be collected from multiple entities, and aggregated at a centralized location, such as a central server. This process of data aggregation not only exposes sensitive information to potential breaches but also conflicts with privacy regulations. Federated learning (FL), a paradigm enabling model training across distributed participants, known as clients, while preserving data privacy and security, effectively addresses the aforementioned challenge of data scarcity and privacy [16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. By allowing data to remain local and only merging model parameters trained with local data, FL effectively enhances DL model performance without compromising data privacy.
However, the practical implementation of FL encounters significant challenges due to data heterogeneity. Variations in production conditions, such as differences in equipment and environmental factors, result in non-independent and identically distributed (non-IID) data across participating clients. For instance, different industrial entities may encounter varying categories of defect samples, leading to discrepancies in the feature representations learned by their models. Such discrepancies make it difficult for different clients to reach a consensus on learned representations. When the uploaded models are aggregated, the global model may become biased towards certain clients, thereby undermining overall performance [23].
@@
To address the aforementioned challenges prevalent in SDC, we propose a personalized FL (PFL) approach, named Adversarial Federated Consensus Learning (AFedCL). It facilitates collaborative training among multiple clients, allowing the development of personalized models tailored to the specific data distributions of each client while preserving privacy. Distinguishing from prior PFL algorithms [26, 27, 28, 29, 30], AFedCL utilizes adversarial training with the global model as a bridge to achieve data distribution alignment across different clients, addressing the problem of global knowledge forgetting in local training explicitly. It significantly improves the ability of local models to extract distribution-invariant features, which are characteristics that remain consistent across varied data distributions among clients, effectively mitigating the adverse effects of data heterogeneity.
@@
    We propose AFedCL, a new PFL method designed to tackle the challenge of data heterogeneity across different clients in SDC within IIoT environments. To the best of our knowledge, this marks the first effort to employ FL in overcoming data scarcity and privacy concerns in SDC.
@@
    We design an adaptive feature fusion module to further improve the efficiency of global knowledge utilization. This module tunes a fusion weight combining global and local features for each client, tailored to their unique levels of global knowledge acquisition, thereby achieving the optimal balance between global and local features.
The proposed method achieves excellent results in extensive experiments on three strip steel SDC datasets. Compared with state-of-the-art methods like FedALA [29], the proposed AFedCL method increases accuracy by up to 5.67%. The rest of this paper is organized as follows. The related work is introduced in Section II, and the FL problem in SDC is formulated in Section III. Section IV presents the proposed method in detail. Numerical experiments are conducted in Section V to verify the effectiveness of the proposed AFedCL framework, and Section VI concludes this paper.
@@ II-A Surface Defect Classification under Data Scarcity
Previous studies have addressed the issue of data scarcity through two main aspects. On one hand, some research focuses on leveraging few-shot learning algorithms. For instance, Kim et al. [15] proposed a siamese neural network with CNN structure to overcome the data scarcity challenge in steel SDC. Similarly, Xiao et al. [31] applied graph embedding and distribution transformation modules, coupled with optimal transport for transductive classification. On the other hand, an alternative strategy involves the use of a generative adversarial network (GAN) to synthesize defect samples, thus enlarging the sample volume. In [32], three GANs are trained to augment the data for surface defects. Furthermore, a cycle-consistent adversarial network with attention mechanism, named AttenCGAN, was proposed to address the challenges of small intra-class differences and data scarcity [33].
@@ II-B Federated Learning for Data Heterogeneity
In light of this, we will introduce local discriminators to dynamically evaluate the efficacy of global knowledge learning and provide real-time feedback to local training. Through adversarial iterative optimization, this methodology enables the distribution alignment among various data distributions, facilitating the dynamic construction of consensus among heterogeneous clients.
@@ III Problem Formulation
2) Data heterogeneity: Variations in production conditions between different industrial entities result in non-IID data. As shown in Fig. 2, different clients hold data with distinct distributions. Since most learning algorithms assume data to be IID, model performance will be degraded due to data heterogeneity.
@@
However, as previously mentioned, the challenge of data heterogeneity severely deteriorates the efficacy of the global model of this overarching one-fit-all paradigm. To address this problem, PFL algorithms shift the focus from a singular, one-fit-all model to the pursuit of distinct models for each participating client. In PFL, the learning objective of each client $k$ could be expressed as
@@ IV-A Overview
As shown in Fig. 3, the overall system consists of a central server and several clients. The entire training process can be summarized into six main steps as follows: :
@@ IV-B Dynamic Consensus Construction
In scenarios where local data distributions vary significantly, models may focus on learning specific characteristics of local data, risking the loss of global knowledge acquired from the global model. To ensure the persistent retention of global knowledge during local training, we develop the first stage of local training (step 2) with ongoing evaluation of global knowledge acquisition and immediate feedback based on adversarial training. Specifically, beyond the classification task, a local discriminator is trained to differentiate between the features extracted by global and local encoders from the identical samples. Concurrently, the local encoder endeavors to confound the discriminator, creating an adversarial game process with the discriminator’s training.
@@
For the discrimination loss $L_{D}^{k,t}$, we still use the cross entropy loss function. However, the approach to calculating discrimination loss differs from that of classification loss in two key ways: firstly, the inputs now involve not just local features but a combination set of local and global features; secondly, the category labels are substituted by binary choices for local and global features. Thus, the feature set for the discrimination tasks, denoted by $F_{k}^{t}$, comprises elements $(f,y)$, where $f$ represents a feature vector either from local encoder $E_{k}^{t}(x)$ with label 00 or global encoder $E_{G}^{t}(x)$ with label $1$, defined as $F_{k}^{t}=\{(E_{k}^{t}(x),0)\}_{(x,y)\in S_{k}}\cup\{(E_{G}^{t}(x),1)\}_{(x,y)\in S_{k}}$. The discrimination loss $L_{D}^{k,t}$ can be calculated by
@@
As communication rounds increase, the discriminator’s ability to distinguish between global and local features strengthens, alongside the encoder’s ability to confuse the discrimination, fostering an adversarial game process. As a result, the similarity between global and local features from identical samples improves. Local encoders imitate the global encoder’s feature patterns, learning to extract discriminative features across different distributions. This alignment process enhances the ability of local encoders to extract distribution-invariant features, which are characteristics that remain consistent across varying distributions, thus constructing a consensus gradually among various clients and enhancing model performance.
@@ V-B3 Data Distribution
The number of clients is set to 5. Data are distributed across two non-IID scenarios to reflect realistic challenges: 1) Disjoint: $c$ classes are randomly allocated to each client, with values of 2, 3, and 4 for the three datasets, respectively; 2) Dirichlet: data is divided using a Dirichlet distribution with $\alpha=0.1$ for all datasets. To simulate the practical scenarios, the number of training samples per client is set to 5, 10, and 20, respectively, across all datasets, reflecting varying degrees of data scarcity. Some crucaial experiment parameters are shown in Table I.
@@ V-B4 Baselines and Evaluation Metrics
We evaluate our method against several state-of-the-art baselines in three categories: 1) training without FL: “No-FL”; 2) one-fit-all FL algorithms: FedAvg [24] and FedProx [25]; 3) PFL algorithms: FedPer [26], FedRep [27], Ditto [28], and FedALA [29]. The evaluation metrics include testing accuracy and F1 score.
@@ V-E Sensitivity Analyses
In this part, we study the performance of the proposed method under various factors: the balanced weight of the discrimination loss function $\lambda$, the backbone of the model, and the number of participating clients $K$.
Firstly, we explore the impact of varying $\lambda$ on the testing performance. Throughout all experiments, $\lambda$ is tuned by selecting the best-performing value from a limited candidate set {0.01, 0.1, 1} for both our method and a competitive baseline, Ditto, which also employs a $\lambda$ to achieve a balance between two loss functions. As shown in Fig. 8, the results reveal that AFedCL consistently outperforms Ditto in testing accuracy and exhibits smaller fluctuation across different $\lambda$ values, demonstrating its stability and superiority.
@@ VI Conclusion
In this paper, we have developed a novel PFL approach, AFedCL, to tackle the challenge of data heterogeneity in SDC. It enables multiple industrial entities in IIoT to collaboratively train personalized models while preserving privacy. Specifically, a dynamic consensus construction strategy has been invented to achieve distribution alignment among different clients through an adversarial game process, thus mitigating performance degradation caused by global knowledge forgetting under non-IID data. Building on this, a consensus-aware aggregation mechanism has been introduced. It assigns aggregating weights based on the global knowledge learning efficacy, measured by each client’s discrimination loss, improving the generalization ability of the global model, and facilitating better consensus construction. Furthermore, an adaptive feature fusion module has been designed. Through the optimization of personalized scalar fusion weights of global and local features, the utilization efficiency of global knowledge is enhanced, thus boosting model performance. The proposed method demonstrates promising results across extensive experiments on three real-world SDC datasets, achieving a notable increase in accuracy of up to 5.67%.

### v2

@@ Abstract
The challenge of data scarcity hinders the application of deep learning in industrial surface defect classification (SDC), as it’s difficult to collect and centralize sufficient training data from various entities in Industrial Internet of Things (IIoT) due to privacy concerns. Federated learning (FL) provides a solution by enabling collaborative global model training across clients while maintaining privacy. However, performance may suffer due to data heterogeneity—discrepancies in data distributions among clients. In this paper, we propose a novel personalized FL (PFL) approach, named Adversarial Federated Consensus Learning (AFedCL), for the challenge of data heterogeneity across different clients in SDC. First, we develop a dynamic consensus construction strategy to mitigate the performance degradation caused by data heterogeneity. Through adversarial training, local models from different clients utilize the global model as a bridge to achieve distribution alignment, alleviating the problem of global knowledge forgetting. Complementing this strategy, we propose a consensus-aware aggregation mechanism. It assigns aggregation weights to different clients based on their efficacy in global knowledge learning, thereby enhancing the global model’s generalization capabilities. Finally, we design an adaptive feature fusion module to further enhance global knowledge utilization efficiency. Personalized fusion weights are gradually adjusted for each client to optimally balance global and local features. Compared with state-of-the-art FL methods like FedALA, the proposed AFedCL method achieves an accuracy increase of up to 5.67% on three SDC datasets.
@@ I Introduction
In recent years, the integration of artificial intelligence with IIoT has provided new automation solutions across various industries, significantly enhancing efficiency and effectiveness[7, 8, 9, 10, 11, 12, 13, 14]. Traditional machine learning classifiers, such as support vector machine (SVM) and random forest [7, 8, 9], have been utilized for automated SDC. These classifiers are typically coupled with some shape-based and texture-based feature extraction methods [10], such as Fourier descriptor and local binary pattern. However, they suffer from limited feature extraction capabilities due to reliance on manual feature engineering, inadequately capturing the diversity of surface defects, thus limiting generalization and adaptability. On the contrary, deep learning (DL)-based methods are capable of learning abstract representations autonomously, without the need for manual feature engineering. In particular, DL-based classifiers built on convolutional neural network (CNN) architectures, such as ResNet and MobileNet, consistently demonstrate superior performance in SDC tasks [11, 12, 13].
Nonetheless, DL-based SDC methods require sufficient data to develop robust models, which a single entity often cannot provide [15]. To overcome this limitation, data must be collected from multiple entities, and aggregated at a centralized location, such as a central server. This process of data aggregation not only exposes sensitive information to potential breaches but also conflicts with privacy regulations. Federated learning (FL), a paradigm enabling model training across distributed participants, known as clients, while preserving data privacy and security, effectively addresses the aforementioned challenge of data scarcity and privacy [16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. By allowing data to remain local and only merging model parameters in the server, FL effectively enhances DL model performance without compromising data privacy.
However, the practical implementation of FL encounters significant challenges due to data heterogeneity. Variations in production conditions, such as differences in equipment and environmental factors, result in non-independent and identically distributed (non-IID) data across participating clients. For instance, different industrial entities may encounter different categories of defect samples, leading to discrepancies in the feature representations learned by their models. Such discrepancies make it difficult for different clients to reach a consensus on learned representations. When the uploaded models are aggregated, the global model may become biased towards certain clients, thereby undermining overall performance [23].
@@
To address the aforementioned challenges, we propose a personalized FL (PFL) approach, named Adversarial Federated Consensus Learning (AFedCL). It facilitates collaborative training among heterogeneous clients, developing personalized models tailored to each client’s specific data distribution while preserving privacy. Distinguishing from prior PFL algorithms [26, 27, 28, 29, 30], AFedCL focuses on the issue of global knowledge forgetting in local training due to distinct data distribution. Through adversarial training, it continuously evaluates the effectiveness of global knowledge learning and updates the model accordingly. Based on this strategy of retaining global knowledge locally for each client, AFedCL aligns the distributions across heterogeneous clients, effectively mitigating performance degradation caused by data heterogeneity.
@@
    We propose AFedCL, a new PFL method designed to tackle the challenge of data heterogeneity across different clients in SDC. To the best of our knowledge, this marks the first effort to employ FL in overcoming data scarcity and privacy concerns in SDC.
@@
    We design an adaptive feature fusion module to further improve the efficiency of global knowledge utilization. This module tunes a fusion weight combining global and local features for each client, thereby achieving the optimal balance between global and local features.
The proposed method achieves superior results in extensive experiments on three strip steel SDC datasets. Compared with state-of-the-art methods like FedALA [29], the proposed AFedCL method increases accuracy by up to 5.67%. The rest of this paper is organized as follows. The related work is introduced in Section II, and the FL problem in SDC is formulated in Section III. Section IV presents the proposed method in detail. Numerical experiments are conducted in Section V to verify the effectiveness of the proposed AFedCL framework, and Section VI concludes this paper.
@@ II-A Surface Defect Classification under Data Scarcity
Previous studies have addressed the issue of data scarcity through two main aspects. On the one hand, some research focuses on leveraging few-shot learning algorithms. For instance, Kim et al. [15] proposed a siamese neural network with CNN structure to overcome the data scarcity challenge in steel SDC. Similarly, Xiao et al. [31] applied graph embedding and distribution transformation modules, coupled with optimal transport for transductive classification. On the other hand, an alternative strategy involves the use of a generative adversarial network (GAN) to synthesize defect samples, thus enlarging the sample volume. In [32], three GANs are trained to augment the data for surface defects. Furthermore, a cycle-consistent adversarial network with attention mechanism, named AttenCGAN, was proposed to address the challenges of small intra-class differences and data scarcity [33].
@@ II-B Federated Learning for Data Heterogeneity
In light of this, we will introduce local discriminators to dynamically evaluate the efficacy of global knowledge learning and provide real-time feedback to local training. Through adversarial iterative optimization, this methodology enables the distribution alignment among various data distributions, facilitating the dynamic construction of consensus among heterogeneous clients to avoid global knowledge forgetting.
@@ III Problem Formulation
2) Data heterogeneity: Variations in production conditions between different entities result in non-IID data. As shown in Fig. 2, different clients hold data with distinct distributions. Since most learning algorithms assume data to be IID, model performance will be degraded due to data heterogeneity.
@@
However, as previously mentioned, the challenge of data heterogeneity severely deteriorates the efficacy of the global model of this overarching one-fit-all paradigm. To address this problem, PFL algorithms shift the focus from a singular, one-fit-all model to the pursuit of distinct models for each participating client. In PFL, the learning objective could be expressed as
@@ IV-A Overview
As shown in Fig. 3, the overall system consists of a central server and several clients. The entire training process can be summarized into six main steps as follows:
@@ IV-B Dynamic Consensus Construction
In scenarios where data distributions vary significantly, models may focus on learning specific characteristics of local data, risking the loss of global knowledge acquired from the global model. To ensure the persistent retention of global knowledge during local training, we develop the first stage of local training (step 2) with ongoing evaluation of global knowledge acquisition and immediate feedback based on adversarial training. Specifically, beyond the classification task, a local discriminator is trained to discriminate between the features extracted by global and local encoders from identical samples. Concurrently, the local encoder endeavors to confuse the discriminator, creating an adversarial game process between the training of the discriminator and the encoder.
@@
For the discrimination loss $L_{D}^{k,t}$, we still use the cross entropy loss function. However, the approach to calculate discrimination loss differs from that of classification loss in two key ways: firstly, the features fed into local discriminator $D_{k}^{t}$ now involve not only local feature $E_{k}^{t}(x)$, but also global feature $E_{G}^{t}(x)$; secondly, the category labels $\{1,2,...,N\}$ are substituted by domain labels $\{0,1\}$, where 00 for label of local features and $1$ for label of global features. Thus, the feature set for the discrimination tasks, denoted by $F_{k}^{t}$ can be defined as $F_{k}^{t}=\{(E_{k}^{t}(x),0)\}_{(x,y)\in S_{k}}\cup\{(E_{G}^{t}(x),1)\}_{(x,y)\in S_{k}}$. The discrimination loss $L_{D}^{k,t}$ can be calculated by
@@
As communication rounds increase, the discriminator’s ability to distinguish between global and local features strengthens, alongside the encoder’s ability to confuse the discrimination. As a result, the similarity between global and local features from identical samples improves. In this way, each client with a distinct data distribution aligns its personalized knowledge with that of the global model. Consequently, the global model serves as a bridge to achieve distribution alignment among heterogeneous clients, dynamically constructing a consensus through several adversarial game processes between the clients and the server.
@@ V-B3 Data Distribution
The number of clients is set to 5. Data are distributed across two non-IID scenarios to reflect realistic challenges: 1) Disjoint: $c$ classes are randomly allocated to each client, with values of 2, 3, and 4 for the three datasets, respectively; 2) Dirichlet: data is divided using a Dirichlet distribution with $\alpha=0.1$ for all datasets. To simulate the practical scenarios, the number of training samples per client is set to 5, 10, and 20, respectively, across all datasets, reflecting varying degrees of data scarcity. Some crucial experiment parameters are shown in Table I.
@@ V-B4 Baselines and Evaluation Metrics
We evaluate the proposed AFedCL method against several state-of-the-art baselines in three categories: 1) training on local private data without FL: “No-FL”; 2) one-fit-all FL algorithms: FedAvg [24] and FedProx [25]; 3) PFL algorithms: FedPer [26], FedRep [27], Ditto [28], and FedALA [29]. The evaluation metrics include testing accuracy and F1 score.
@@ V-E Sensitivity Analyses
Here, we study the robustness of the proposed method under various factors: the balanced weight of the discrimination loss function $\lambda$, the backbone of the model, and the number of participating clients $K$.
Firstly, we explore the impact of varying $\lambda$ on the testing performance. Throughout all experiments, $\lambda$ is tuned by selecting the best-performing value from a candidate set $\{0.01,0.1,1\}$ for both our method and a competitive baseline, Ditto, which also employs a $\lambda$ to achieve a balance between two loss functions. As shown in Fig. 8, the results reveal that AFedCL consistently outperforms Ditto in testing accuracy and exhibits smaller fluctuation across different $\lambda$ values, demonstrating its stability and superiority.
@@ VI Conclusion
In this paper, we have developed a novel PFL approach, AFedCL, to tackle the challenge of data heterogeneity in SDC. It enables multiple industrial entities in IIoT to collaboratively train personalized models while preserving privacy. Specifically, a dynamic consensus construction strategy has been invented to achieve distribution alignment among different clients through an adversarial game process, thus mitigating performance degradation caused by global knowledge forgetting under non-IID data. Building on this, a consensus-aware aggregation mechanism has been introduced. It assigns aggregating weights based on the global knowledge learning efficacy, measured by each client’s discrimination loss, improving the generalization ability of the global model. Furthermore, an adaptive feature fusion module has been designed. Through the optimization of personalized scalar fusion weights of global and local features, the utilization efficiency of global knowledge is enhanced, thus boosting model performance. The proposed method demonstrates promising results across extensive experiments on three real-world SDC datasets, achieving a notable increase in accuracy of up to 5.67%.